# Holographic Intelligence Surface Assisted Integrated Sensing and Communication

### v1

@@ III-C Problem Formulation in Discrete Domain
Due to the non-convex max-min objective function and coupled optimized variables in objective and constraints, (20) is non-convex and challenging to solve. To circumvent the difficulties, we develop an efficient AO algorithm to address the joint transmit-receive beamforming in the HISAC system. Specifically, the ABS is proposed to accelerate the convergence of the SDR-based bisection search in the transmit beamforming, and a Rayleigh quotient-based method is developed for the receive beamformer design.
IV Joint Transmit-Receive Beamforming in HISAC System
We begin with introducing the transmit HIS beamforming in Section IV-A, which we then utilize to design the receive beamforming with fixed transmit beamforming in Section IV-B.
@@ IV-A1 SDR-Based Algorithm
$\eta_{c}(\boldsymbol{R};k)=\frac{\boldsymbol{f}_{k}^{H}\boldsymbol{R}_{k}\boldsymbol{f}_{k}}{\boldsymbol{f}_{k}^{H}\left(\boldsymbol{R}-\boldsymbol{R}_{k}\right)\boldsymbol{f}_{k}+\frac{\sigma_{c}^{2}}{\kappa^{2}Z^{2}_{0}}},~{}k=1,...,K,$ (22)
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}l=1,...,M,$
where $\boldsymbol{R}_{k}=\boldsymbol{w}_{k}\boldsymbol{w}_{k}^{H}$ and $\boldsymbol{R}=\boldsymbol{W}\boldsymbol{W}^{H}$. With the given receive beamformer $\boldsymbol{Q}=\left[\boldsymbol{q}_{1},...,\boldsymbol{q}_{M}\right]$, we propose to maximize the minimal sensing SINR by optimizing the total covariance matrix $\boldsymbol{R}$ and its sub-element $\boldsymbol{R}_{k}$. Then, the transmit beamforming design problem is rewritten as
$\centering\max_{\boldsymbol{R},\boldsymbol{R}_{1},...,\boldsymbol{R}_{K}}\min_{l=1,...,M}\ \@add@centering$ $\eta_{r}(\boldsymbol{R};l),$ (23a)
${\rm s.t.}$ $\eta_{c}(\boldsymbol{R};k)\geq\Gamma_{c},~{}k=1,...,K,$ (23b)
$\boldsymbol{R}\succeq 0,$ (23c)
$\boldsymbol{R}_{k}\succeq 0,~{}k=1,...,K,$ (23d)
$\boldsymbol{R}-\sum_{k=1}^{K}\boldsymbol{R}_{k}\succeq 0,$ (23e)
${\rm tr}(\boldsymbol{R})\leq P_{T},$ (23f)
${\rm rank}(\boldsymbol{R}_{k})=1,~{}k=1,...,K.$ (23g)
The transmit beamforming design problem outlined above is still non-convex. To deal with it, we first omit the rank-one constraints and transform (23) into a series of feasibility-checking [39] sub-problems by SDR as
@@
7:  $\boldsymbol{R},\boldsymbol{R}_{1},...,\boldsymbol{R}_{K}$.
Suppose the optimal solution $\boldsymbol{\hat{R}}_{1},...,\boldsymbol{\hat{R}}_{K}$ of (24) are exactly rank-one. In that case, the relaxation used in SDR is tight, i.e., the solution to (24) coincides with the original non-convex problem (23). If the rank-one requirement is not fulfilled, the following proposition enables us to reconstruct the rank-one solution for the original problem and generate the communication beamformer $\boldsymbol{\hat{W}}_{c}=[\boldsymbol{\hat{w}}_{1},...,\boldsymbol{\hat{w}}_{K}]$ and sensing beamformer $\boldsymbol{\hat{W}}_{r}$, respectively.
Proposition 2
@@
$\centering\boldsymbol{\hat{w}}_{k}=\left(\boldsymbol{f}_{k}^{H}\boldsymbol{\hat{R}}_{k}\boldsymbol{f}_{k}\right)^{-\frac{1}{2}}\boldsymbol{\hat{R}}_{k}\boldsymbol{f}_{k},~{}k=1,...,K,$ (27)
@@
The optimal solution $\boldsymbol{\bar{R}}_{1},...,\boldsymbol{\bar{R}}_{K}$ for (20) is reconstructed as $\boldsymbol{\bar{R}}_{k}=\boldsymbol{\hat{w}}_{k}\boldsymbol{\hat{w}}_{k}^{H}$, for $k=1,...,K$, which is exactly rank-one and optimal.
Proof:
See Appendix -B. ∎
@@ IV-A2 Adaptive Bisection Searching Method
0:  $\Gamma_{c}$. Initial point: $\Gamma_{r,start}=\frac{P_{T}A_{T}}{\sigma_{R}^{2}}-K\Gamma_{c}$.
1:  while $1$ do
@@
5:  $\Gamma_{r,end}\leftarrow\Gamma_{r,start}+\Gamma_{c}$.
6:  while $1$ do
@@
10:  $\Gamma_{r,start}\leftarrow\Gamma_{r,end}-\Gamma_{c}$.
10:  $\left[\Gamma_{r,start},\Gamma_{r,end}\right]$.
@@
Finally, combined with a shrinked searching range obtained by the proposed ABS approach, the SDR-based bisection method in Algorithm 1 is employed to calculate the solution to the problem (24), with $t$ serving as the solution indicator. Once the search range meets $\Gamma_{r,end}-\Gamma_{r,start}\leq\epsilon_{1}$ and $t\geq 0$, the iterative algorithm ends to obtain the optimal solution $\boldsymbol{R},\boldsymbol{R}_{1},...,\boldsymbol{R}_{K}$, which we then utilize to calculate the transmit beamformer $\boldsymbol{W}$ via (27) and (28).
@@ IV-B HIS Receive Beamforming with Fixed Transmit Beamforming
10:  $\boldsymbol{W}$, $\boldsymbol{Q}$.
Specifically, the transmit beamformers and receive beamformers are updated alternately until a given convergence tolerance between two iterative steps is met, i.e., $\left|\Gamma_{r}^{*(i)}-\Gamma_{r}^{*(i-1)}\right|<\epsilon_{2}$, where $\Gamma_{r}^{*(i)}$ and $\Gamma_{r}^{*(i-1)}$ denote the minimal sensing SINR in $(i)$-the and $(i-1)$-th iterations, respectively.
@@ V-A Simulation Setup
Without specified otherwise, the simulation parameters are summarized in Table. I, and the geometry of the simulation scenario is illustrated in Fig. 3.
@@
The $\mathcal{P}_{S}$ and $\mathcal{P}_{R}$ are located in the XOY plane adjacently. Simultaneously, all communication users and radar targets are located on the far-field radiation region of the HIS, represented by polar coordinates. Additionally, to underscore the superiority of the proposed HISAC system, we utilize the discrete array counterpart, where antennas are deployed with half-wavelength spacing, as a baseline.
TABLE I: The basic simulation parameters
@@
Position of the target $1$ $(30^{\circ},90^{\circ})$
Position of the target $2$ $(30^{\circ},45^{\circ})$
@@
Compared to the HIS, we assume the discrete-array-based system consists of discrete patches (isotropic source) with half-wavelength inter-distance. The aperture size of the discrete array remains the same as the HIS, allowing $D=DxDy=\frac{L_{x}}{\lambda/2}\frac{L_{y}}{\lambda/2}$ antennas to be optimized. According to [43], the efficient aperture size of the isotropic source is given by $\frac{\lambda^{2}}{4\pi}$. Therefore, the maximal radiation power of the discrete array with $D$ elements can be calculated as $\frac{P_{T}A_{T}}{\pi}.$ Based on this setup, compared to the HISAC system, the round-trip performance degradation of the discrete array counterpart is $\frac{1}{\pi^{2}}$.
@@ V-B1 Algorithm Convergence
As can be seen, the indicator $t$ converges to 00, implying the optimal solution to (24) is found. Specifically, when $t>0$, the potential value $\Gamma_{r}$ is feasible but not optimal; when $t<0$, the potential value $\Gamma_{r}$ is infeasible. If and only if $t\to 0$, the solution of (24) converges to the optimal value. Therefore, the minimal sensing SINR oscillates and converges to the maximal value after multiple iterations.
@@ V-B2 Multi-User Single-Target Scenario
In Fig.  5, we evaluate the sensing SINR versus the maximal transmit power $P_{T}$ for different ISAC configurations. We observe that our proposed HISAC system outperforms the discrete array counterpart in the considered range of power budget.
@@
As can be seen, the radar performance improves with the increase of the aperture size, and the sensing SINR of our proposed HISAC system exhibits a progressive enhancement as the Fourier expansion order $N$ increases outperforming its discrete array counterpart even when $N$ is moderately large. However, if the Fourier expansion order is too small and insufficient to capture the HIS radiation characteristics, the radar performance will degrade severely at the large aperture size region, potentially becoming worse than its discrete array counterpart.
In Fig. 7 and Fig. 7, we illustrate the transmit and receive beampatterns of the HIS and that of the discrete array at $\theta=30^{\circ}$ plane, respectively. Based on these normalized beampatterns, we study the beampattern gain between the HIS-based and discrete-array-based system. As can be seen, in the transmit phase, as the Fourier expansion order $N$ increases, the beamwidth of the HIS’s transmit beampattern tightens, where the peak power reaches its maximum until the $N$ is sufficient for characterizing the HIS radiation. Meanwhile, the peak value of the transmit beampattern based on the discrete array demonstrates $4.9$ dB degradation compared to HIS, which agrees with $\pi$. In addition, for the receive beampattern w.r.t. target $1$, the peak value of the HIS’s beampattern outperforms that of the discrete array with a $9.7$ dB enhancement. As a result, the sensing SINR for target $1$ can be significantly improved after the receive beamforming under the proposed HISAC transceiver architecture. The reason is that both the HIS’s transmit and receive beamformer are optimized at the receiver side, leading to the overall sensing SINR improvement in our proposed architecture.
Figure 7: HIS beampatterns for target $1$. (a) Transmit beampattern. (b) Receive beampattern.
In Fig. 8, we explore the sensing SINR under varying communication SINR requirement. Specifically, we evaluate sensing SINR versus different communication threshold $\Gamma_{c}$. The physical aperture size is $A_{T}=0.6\times 0.6$ m2 to demonstrate the superior sensing SINR difference between Fourier expansion order $N$ and its discrete array counterpart. A clear trade-off between sensing and communication is presented, in which the sensing SINR decreases as the communication SINR requirement increases. In addition, we observe that our proposed HISAC system shows robustness compared to the discrete array system, possessing more significant sensing performance gain with higher communication SINR threshold. Specifically, as $\Gamma_{c}$ increases, the sensing SINR gain of the HISAC system over the discrete array-based ISAC system increases from $7.4$ dB to $15.6$ dB, demonstrating the advantages of incorporating HIS into ISAC systems.
@@ V-B3 Multi-User Multi-Target Scenario
An additional radar target positioned at $(30^{\circ},45^{\circ})$ is introduced in the multi-user multi-target scenario. In Fig. 9, we compare the minimal sensing SINR among radar targets for different Fourier expansion order $N$ versus $P_{T}$. As observed, due to inter-target interference, the sensing SINR undergoes a $3$ dB degradation, compared to the interference-free upper bound. Nevertheless, our proposed HISAC system still outperforms its discrete array counterpart with $9.7$ dB sensing SINR enhancement. The insight here is that the number of radar targets will not affect the sensing SINR enhancement between the HIS and the discrete array. This is because the radiation power of both architectures is independent of the number/directions of targets. Therefore, when the Fourier expansion order $N$ is sufficient, our system can still achieve power gain around $\pi^{2}$.
@@
In Fig. 10, we illustrate the normalized transmit and receive beampatterns at $\theta=30^{\circ}$ plane to demonstrate the impact of multi-target sensing on the minimal sensing SINR. As observed in Fig. 10, the main lobe of the HIS transmit beampattern is split into two main lobes with the same peak level. Compared to the discrete array, the transmit beampattern of HIS exhibits a $4.9$ dB enhancement in each direction of targets. As can be seen in Fig. 10, the receiving beamforming strengthens the beampattern’s peak values in the directions of targets while suppressing those of the clutters (targets located in other directions) deeply. Specifically, for the receive beamforming w.r.t. target $1$, the direction w.r.t. target $1$ is enhanced while the direction w.r.t. target $2$ is suppressed and vice versa. Note that the peak value of HIS still maintains an enhancement around $9.7$ dB compared to the discrete array, regardless of the directions of the receive beamforming. Therefore, our proposed algorithm is efficient in addressing the multi-user multi-target scenario and indicates that the sensing SINR improvement is attributed to the efficient aperture and is independent of the directions of targets.
VI Conclusions

### v2

@@ III-D Problem Formulation in Discrete Domain
Due to the non-convex max-min objective function and coupled optimized variables in objective and constraints, (24) is non-convex and challenging to solve. To circumvent the difficulties, we develop an efficient AO algorithm to address the joint transmit-receive beamforming in the HISAC system.
IV Joint Transmit-Receive Beamforming in HISAC
In this section, we develop an AO algorithm to address the joint transmit-receive beamforming in our proposed HISAC system under fully-digital architecture. We begin with introducing the transmit HIS beamforming in Section IV-A, which we then utilize to design the receive beamforming with fixed transmit beamforming in Section IV-B. In addition, we analyze the complexity and convergence of the proposed algorithm in Section IV-C.
@@ IV-A1 SDR-Based Algorithm
$\eta_{c}(\boldsymbol{R};k)=\frac{\boldsymbol{f}_{k}^{H}\boldsymbol{R}_{k}\boldsymbol{f}_{k}}{\boldsymbol{f}_{k}^{H}\left(\boldsymbol{R}-\boldsymbol{R}_{k}\right)\boldsymbol{f}_{k}+\frac{\sigma_{c}^{2}}{\kappa^{2}Z^{2}_{0}}},~k=1,.,K,$ (26)
@@
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~l=1,.,M,$
where $\boldsymbol{R}_{k}=\boldsymbol{w}_{k}\boldsymbol{w}_{k}^{H}$ and $\boldsymbol{R}=\boldsymbol{W}\boldsymbol{W}^{H}$. With the given receive beamformer $\boldsymbol{Q}=\left[\boldsymbol{q}_{1},...,\boldsymbol{q}_{M}\right]$, we propose to maximize the minimal sensing SINR by optimizing the total covariance matrix $\boldsymbol{R}$ and its sub-element $\boldsymbol{R}_{k}$.
To deal with it, we first omit the rank-one constraints and transform (24) into a series of feasibility-checking [40] sub-problems by SDR as
@@
[…]
@@
$\centering\boldsymbol{\hat{w}}_{k}=\left(\boldsymbol{f}_{k}^{H}\boldsymbol{\hat{R}}_{k}\boldsymbol{f}_{k}\right)^{-\frac{1}{2}}\boldsymbol{\hat{R}}_{k}\boldsymbol{f}_{k},~k=1,.,K,$ (30)
@@
The optimal solution $\boldsymbol{\bar{R}}_{1},...,\boldsymbol{\bar{R}}_{K}$ for (24) is reconstructed as $\boldsymbol{\bar{R}}_{k}=\boldsymbol{\hat{w}}_{k}\boldsymbol{\hat{w}}_{k}^{H}$, for k=1,,Kk=1,...,K, which is exactly rank-one and optimal.
@@ IV-A2 Adaptive Bisection Searching Method
[…]
1:while 11 do
@@
[…]
6:while 11 do
@@
[…]
10:[$\left[\Gamma_{r,start},\Gamma_{r,end}\right]$.
@@
Finally, combined with a confined searching range obtained by the proposed ABS approach, the SDR-based bisection method in Algorithm 1 is employed to calculate the solution to the problem (27), with $t$ serving as the solution indicator. Once the search range meets $\Gamma_{r,end}-\Gamma_{r,start}\leq\epsilon_{1}$ and $t\geq 0$, the iterative algorithm ends to obtain the optimal solution $\boldsymbol{R},\boldsymbol{R}_{1},...,\boldsymbol{R}_{K}$, which we then utilize to calculate the transmit beamformer $\boldsymbol{W}$ via (30) and (31).
@@ IV-B HIS Receive Beamforming with Fixed Transmit Beamforming
[…]
Specifically, the transmit beamformers and receive beamformers are updated alternately until a given convergence tolerance between two iterative steps is met, i.e., $\left|\Gamma_{r}^{*(i)}-\Gamma_{r}^{*(i-1)}\right|<\epsilon_{2}$, where $\Gamma_{r}^{*(i)}$ and $\Gamma_{r}^{*(i-1)}$ denote the minimal sensing SINR in $(i)$-the and (i1)(i-1)-th iterations, respectively.
IV-C Complexity and Convergence Analysis
IV-C1 Computational Complexity Analysis
The computational complexity of the proposed AO algorithm comprises two parts: transmit-side optimization and receive-side update. The transmit-side optimization is formulated as an SDP, whose complexity dominates the overall computation. The receive-side update reduces to a generalized Rayleigh quotient problem and admits a closed-form solution. Consequently, each iteration of the proposed algorithm involves solving an SDP and then performing a closed-form update. In the following, we provide the computational complexity of each sub-algorithm.
  • Optimization of transmit beamforming matrices $\boldsymbol{W}_{c}$, $\boldsymbol{w}_{r}$ with fixed receive HIS configuration: The SDP problem in (26) is solved by the ABS-SDR approach. Specifically, given the potential values of $\Gamma_{r}$, the number of execution steps of the ABS $I_{o}$. Given the solution accuracy $\epsilon_{1}$, the complexity of using the SDR to solve (26) is $\mathcal{O}\left(N^{3.5}\log(1/\epsilon_{1})\right)$ [44]. Therefore, the computational complexity of transmit beamforming algorithm is $I_{o}\times\mathcal{O}\left(N^{3.5}\log(1/\epsilon_{1})\right)$.
  • Optimization of receive HIS configuration with fixed transmit beamforming matrices: Since we suggest the generalized Rayleigh quotient algorithm to address (35), the complexity of this algorithm is $\mathcal{O}\left(N^{3}\right)$.
As a result, the computational complexity of the proposed alternating algorithm for each iteration is $I_{o}\times\mathcal{O}\left(N^{3.5}\log(1/\epsilon_{1})\right)+\mathcal{O}\left(N^{3}\right)$.
IV-C2 Convergence Analysis
Due to the non-convex nature of the joint transmit-receive beamforming problem, global optimality of Algorithm 3 cannot be formally guaranteed. However, the proposed AO algorithm ensures convergence to a stationary point, as each subproblem yields either a convex formulation or a closed-form solution, and guarantees that the objective value does not decrease in each iteration. On the receiver side, the HIS configuration is optimized using a generalized Rayleigh quotient, which yields a closed-form solution without iterative updates. Therefore, convergence analysis primarily focuses on the transmit-side HIS optimization and the overall AO process. For the transmit-side configuration, we adopt an adaptive algorithm based on the SDR framework, termed ABS-SDR. The algorithm iteratively adjusts the search interval $\left[\Gamma_{r,start},\Gamma_{r,end}\right]$ according to the achieved sensing SINR. If the SINR improves, the new configuration is accepted and the search region is tightened. Otherwise, the algorithm reverts to the previous configuration. Since both sensing and communication SINR are inherently bounded due to physical limitations such as total transmit power, this ascent property guarantees the convergence of the ABS-SDR algorithm. As both subproblems exhibit the ascent property and are constrained within a limited power constraint, the overall AO algorithm is guaranteed to converge.
@@ V-A Simulation Setup
Without specified otherwise, the simulation parameters are summarized in Table II, and the geometry of the simulation scenario is illustrated in Fig. 4.
@@
The $\mathcal{P}_{S}$ and $\mathcal{P}_{R}$ are located in the XOY plane adjacently. Simultaneously, all communication users and radar targets are located on the far-field radiation region of the HIS, represented by polar coordinates. Additionally, to underscore the superiority of the proposed HISAC system, we utilize the discrete array counterpart, where antennas are deployed with half-wavelength spacing, as a baseline. Note that the transmit power $P_{T}$ is computed based on the continuous-aperture current model, where the unit is expressed in mA2 rather than in watt [16]. Moreover, the truncation order is not optimized on a per-scenario basis. Instead, it is determined by the physical aperture size and the carrier wavelength, which jointly limit the wavenumber bandwidth of the radiated field. Once the aperture is fixed, the truncation order can be selected according to (16), and remains unchanged across various communication and sensing scenarios.
TABLE II: The basic simulation parameters
@@
Position of the target 11 $(30^{\circ},90^{\circ})$
Position of the target 22 $(30^{\circ},45^{\circ})$
@@
Compared to the HIS, we assume the discrete-array-based system consists of discrete patches (isotropic source) with half-wavelength inter-distance. The aperture size of the discrete array remains the same as the HIS, allowing $D=DxDy=\frac{L_{x}}{\lambda/2}\frac{L_{y}}{\lambda/2}$ antennas to be optimized. According to [45], the efficient aperture size of the isotropic source is given by $e_{a}=\frac{\lambda^{2}}{4\pi}$. Therefore, the maximal radiation power of the discrete array with $D$ elements can be calculated as $\frac{P_{T}A_{T}}{\pi}.$ Based on this setup, compared to the HISAC system, the round-trip performance degradation of the discrete array counterpart is $\left(\frac{e_{a}D}{L_{x}L_{y}}\right)^{2}=\frac{1}{\pi^{2}}$.
@@ V-B1 Algorithm Convergence
As can be seen, the indicator $t$ converges to 0, implying the optimal solution to (27) is found. Specifically, when $t>0$, the potential value $\Gamma_{r}$ is feasible but not optimal; when $t<0$, the potential value $\Gamma_{r}$ is infeasible. If and only if $t\to 0$, the solution of (27) converges to the optimal value. Therefore, the minimal sensing SINR oscillates and converges to the maximal value after multiple iterations.
@@ V-B2 Multi-User Single-Target Scenario
In Fig.  6, we evaluate the sensing SINR versus the maximal transmit power $P_{T}$ for different ISAC configurations. We observe that our proposed HISAC system outperforms the discrete array counterpart in the considered range of power budget.
@@
As can be seen, the radar performance improves with the increase of the aperture size, and the sensing SINR of our proposed HISAC system exhibits a progressive enhancement as the Fourier expansion order $N$ increases, outperforming its discrete array counterpart even when $N$ is moderately large. However, if the Fourier expansion order is too small to adequately capture the HIS radiation characteristics, the radar performance degrades significantly in the large-aperture region and may fall below that of the discrete array counterpart.
In Fig. 8 and Fig. 8, we illustrate the transmit and receive beampatterns of the HIS and that of the discrete array at $\theta=30^{\circ}$ plane, respectively. Based on these normalized beampatterns, we study the beampattern gain between the HIS-based and discrete-array-based system. As can be seen, in the transmit phase, as the Fourier expansion order $N$ increases, the beamwidth of the HIS’s transmit beampattern tightens, where the peak power reaches its maximum until the $N$ is sufficient for characterizing the HIS radiation. Meanwhile, the peak value of the transmit beampattern based on the discrete array demonstrates $4.9$ dB degradation compared to HIS, which agrees with $\pi$. In addition, for the receive beampattern w.r.t. target 11, the peak value of the HIS’s beampattern outperforms that of the discrete array with a $9.7$ dB enhancement. As a result, the sensing SINR for target $1$ can be significantly improved after the receive beamforming under the proposed HISAC transceiver architecture. The reason is that both the HIS’s transmit and receive beamformers are optimized at the receiver side, leading to the overall sensing SINR improvement in our proposed architecture.
Figure 8: HIS beampatterns for target 11. (a) Transmit beampattern. (b) Receive beampattern.
In Fig. 9, we explore the sensing SINR under varying communication SINR requirements. Specifically, we evaluate sensing SINR versus different communication threshold $\Gamma_{c}$. The physical aperture size is $A_{T}=0.6\times 0.6$ m2 to demonstrate the superior sensing SINR difference between Fourier expansion order $N$ and its discrete array counterpart. A clear trade-off between sensing and communication is presented, in which the sensing SINR decreases as the communication SINR requirement increases. In addition, we observe that our proposed HISAC system shows robustness compared to the discrete array system, possessing more significant sensing performance gain with higher communication SINR threshold. Specifically, as $\Gamma_{c}$ increases, the sensing SINR gain of the HISAC system over the discrete array-based ISAC system increases from $7.4$ dB to $15.6$ dB, demonstrating the advantages of incorporating HIS into ISAC systems.
@@ V-B3 Multi-User Multi-Target Scenario
To evaluate the scalability and effectiveness of the proposed HISAC framework, we provide comprehensive simulations under a multi-user and multi-target scenario. An additional radar target positioned at $(30^{\circ},45^{\circ})$ is introduced in the multi-user multi-target scenario. In Fig. 10, we compare the minimal sensing SINR among radar targets for different Fourier expansion order $N$ versus $P_{T}$. As observed, due to inter-target interference, the sensing SINR undergoes a $3$ dB degradation, compared to the interference-free upper bound. Nevertheless, our proposed HISAC system still outperforms its discrete array counterpart with $9.7$ dB sensing SINR enhancement. The insight here is that the number of radar targets will not affect the sensing SINR enhancement between the HIS and the discrete array. This is because the radiation power of both architectures is independent of the number/directions of targets. Therefore, when the Fourier expansion order $N$ is sufficient, our system can still achieve power gain around $\pi^{2}$.
@@
In Fig. 11, we illustrate the normalized transmit and receive beampatterns at $\theta=30^{\circ}$ plane to demonstrate the impact of multi-target sensing on the minimal sensing SINR. As observed in Fig. 11, the main lobe of the HIS transmit beampattern is split into two main lobes with the same peak level. Compared to the discrete array, the transmit beampattern of HIS exhibits a $4.9$ dB enhancement in each direction of targets. As can be seen in Fig. 11, the receiving beamforming strengthens the beampattern’s peak values in the directions of targets while suppressing those of the clutters (targets located in other directions) deeply. Specifically, for the receive beamforming w.r.t. target 11, the direction w.r.t. target $1$ is enhanced while the direction w.r.t. target $2$ is suppressed and vice versa. Note that the peak value of HIS still maintains an enhancement around $9.7$ dB compared to the discrete array, regardless of the directions of the receive beamforming. Therefore, our proposed algorithm is efficient in addressing the multi-user multi-target scenario and indicates that the sensing SINR improvement is attributed to the efficient aperture and is independent of the directions of targets.
Figure 12: SINRs for sensing and communication versus imperfect CSI bias.
To further evaluate the robustness of the proposed HISAC system under channel uncertainty, we conduct simulations under imperfect CSI conditions. The angular deviation $\Delta_{\theta}$ represents the mismatch between the estimated and actual angle of arrival/departure for User 1, thereby capturing CSI inaccuracy. We compare the sensing and communication SINR across all radar targets and communication users for both HIS-based and discrete-array-based systems. As shown in Fig. 12, the results reveal a distinct trade-off between the Fourier truncation order $N$ of the HIS and its robustness to CSI error. Specifically, higher-order HIS configurations achieve better radiation fidelity and stronger beamforming gain under ideal CSI, but become more sensitive to imperfect CSI. Nevertheless, all HIS configurations consistently outperform their discrete-array-based counterparts, even under moderate CSI error. Results demonstrate that the efficient aperture utilization of HIS enables stable SINR performance for both sensing and communication tasks.
VI Conclusion and Discussion